OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 5 y + 3 = 2x 2x + 2x - 3 = 5 2x + 3 - 2x = 5 2x + 2x + 3 = 5

OpenStudy (texaschic101):

y + 3 = 2x y = 2x - 3 now sub 2x - 3 in for y back into the first equation 2x + y = 5 2x + (2x - 3) = 5

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thank you so much! =))